# approximate nearest neighbour IOU intersection over union

### v1

@@ Abstract
Decomposing 3D assets into material parts is a common task for artists and creators, yet remains a highly manual process. In this work, we introduce Select Any Material (SAMa), a material selection approach for various 3D representations. Building on the recently introduced SAM2 video selection model, we extend its capabilities to the material domain. We leverage the model’s cross-view consistency to create a 3D-consistent intermediate material-similarity representation in the form of a point cloud from a sparse set of views. Nearest-neighbor lookups in this similarity cloud allow us to efficiently reconstruct accurate continuous selection masks over objects’ surfaces that can be inspected from any view. Our method is multiview-consistent by design, alleviating the need for contrastive learning or feature-field pre-processing, and performs optimization-free selection in seconds. Our approach works on arbitrary 3D representations and outperforms several strong baselines in terms of selection accuracy and multiview consistency. It enables several compelling applications, such as replacing the diffuse-textured materials on a text-to-3D output, or selecting and editing materials on NeRFs and 3D-Gaussians. Webpage: https://mfischer-ucl.github.io/sama.
@@ 1 Introduction
Understanding the materials around us is an extremely common task for humans, but remains challenging for machine vision approaches. In this paper we focus on material selection on 3D objects.
Existing work on material understanding has mostly focused on the 2D image domain, addressing tasks like segmentation [58, 41, 60, 4], reconstruction [14, 15, 57, 26], generation [63, 53, 62] or, more recently, material selection [56]. In material selection, the task is to select all image pixels that share the same material, given a user prompt (typically a click on the material of interest), regardless of variations in shading or object boundaries. This is an important distinction from semantic or object selection, which aim to identify individual objects or semantic object groups. For instance, when clicking on a wood material, we neither want just the current object that the material is applied on (e.g., a single chair), nor the semantic entity or group that the material belongs to (e.g., all chairs). Instead, we want to return all parts made of that specific type of wood. We differentiate selection and segmentation, where the former aims to identify one material based on a user click prompt, while the latter targets a complete decomposition into different material regions. We follow the material definition of established works [56, 13] and consider two materials similar if they share the same texture and reflectance properties.
@@
Material selection becomes especially relevant in the light of recent generative 3D asset creation and image/text-to-3D workflows. Current methods either provide non-parametric implicit representations (e.g., Neural Radiance Fields) or unstructured output (as in triangle soups and baked textures produced by image/text-to-mesh methods [25, 38, 21, 45]), both of which are challenging to use for artists and downstream tasks. Material selection, in this context, has a wide range of downstream applications, e.g., enhancing the X-to-3D workflow with material masks, improving the editability of 3D reconstructions (e.g., through material replacement), or extracting areas of similar materials as a prior for observation sharing in inverse rendering.
@@
However, most models targeting material-related tasks, including selection, do not trivially extend to the 3D domain, as they are trained on 2D images and therefore have no incentive for producing multiview-consistent predictions [17]. Moreover, the 3D domain contains inherent challenges and ambiguities like self- or dis-occlusions and view-dependent effects, and requires accurate propagation of the model predictions into novel, unseen views. Recent research has therefore developed algorithms to address the problems from multiview-inconsistent predictions that arise when lifting 2D (object) selection to 3D, predominantly via pre-processing noise-consolidation steps such as feature-field distillation [32] or contrastive (similarity) learning [30], both of which are time-consuming.
@@
In this work, we close the gap between material selection in 2D and 3D and propose an efficient and accurate material selection and segmentation method for 3D objects. One of our core insights is that we can draw parallels between video selection and accurate 3D selection, since in both video and 3D, the selected elements have to be consistent across frames (or views), regardless of object and camera movement or differences in shading and occlusions. We thus propose to re-purpose SAM2’s recent progress in object selection across video [49] for materials. We achieve this by fine-tuning parts of the model for (video) material selection on a custom-made video dataset with dense per-pixel, per-frame annotations. We show that the use of videos for fine-tuning is key to achieving high quality in 3D.
Our approach is inherently multiview-consistent thanks to its video training. Therefore, once our SAMa is trained, it enables selection in less than two seconds from the initial click on an arbitrary 3D object. Selection visualization from different views can then be performed in under 10 milliseconds, enabling interactive visualization and editing of our selection results.
@@
Importantly, our approach supports selection on any 3D representation that can be rendered to an image and queried for depth. We show selection results on meshes, radiance fields and 3D Gaussians. We evaluate our method qualitatively and quantitatively, both in terms of selection quality and 3D consistency, and show that it improves significantly over existing work and several strong baselines. Finally, we demonstrate multiple applications such as object segmentation into material IDs and NeRF/Gaussian editing.
@@
  • Adaptation of a video-object-selection model to material selection on 3D shapes, by training on a novel rendered video dataset.
  • Fast and efficient 3D projection, enabled by cross-frame consistency.
@@
Throughout this paper, we will show the user-provided input clicks with respect to which we select materials as black circles. We will show the material similarity to these clicks in false colors, with blue and red indicating low and high, respectively.
@@ 2 Related work
Material selection in 2D.
Most prior works in material segmentation rely on hand-crafted features [51, 5, 40, 24, 48] or focus on images of flat surfaces [34, 26, 35, 44, 11]. Recently, Sharma et al. [56] proposed Materialistic, a model based on DINO-ViT [7] features, trained to predict the material similarity between a query pixel and all other pixels in a natural image. We find that Materialistic struggles with accurate material selection on 3D objects for two reasons: (1) it is trained for full-scene photographs, leading to limited selection precision on objects, and (2) its selections are not sufficiently consistent to be lifted to varying 3D views.
Closely related, the Segment Anything Model (SAM) [31] uses a ViT trained to predict similarity between pixels. As its training data is object-selection specific and not material-aware, SAM requires many separate clicks to perform even moderately well on materials. The more recent SAM2 [49] also targets object selection, but introduces support for temporally coherent predictions across video frames. We find that neither SAM nor SAM2 perform well on material selection, except in the special case of an object made of a single material. However, once fine-tuned for materials, SAM2’s cross-frame selection consistency enables our fast selection lifting to 3D.
@@
Due to the scarcity of annotated 3D data and the increased computational complexity compared to 2D, many approaches have attempted to lift available 2D predictions from multiple views to a shared 3D representation. We here discuss approaches that tackle selection and segmentation.
@@ 3 Method
Our approach targets material selection on 3D representations. Existing methods focus mostly on selection in 2D images [56], and their extension to 3D is not trivial as the underlying vision models are not consistent across views. They also do not cope well with selected material regions going out of frame or changes in the frames’ background.
Instead of enforcing 3D consistency through per-asset optimization and feature consolidation [32, 30, 23, 66], we take inspiration from memory priors in recent video models [49, 61] which show good cross-frame consistency. Since renders of a 3D object from a smooth camera trajectory are not markedly different from a video, we propose to adapt SAM2 [49] to material selection by fine-tuning it, including its memory bank components, on a material-specific video dataset that we design. Once fine-tuned, given an image of a rendered object and a (clicked) pixel, the model outputs a floating-point map that encodes the similarity between the clicked pixel’s material and all other pixels. Thresholding such a similarity map yields a binary selection mask.
However, fine-tuning alone does not yet lead to interactive selection in 3D. While it enables material selection from novel views with good consistency, it is not yet efficient, as it requires querying the model for each such view. Furthermore, cross-frame consistency can still exhibit artifacts in challenging cases and long frame sequences. To alleviate inconsistencies, we consolidate the 2D similarity maps of a sparse set of key-frames into a 3D similarity point cloud. Using this point cloud and nearest-neighbor queries, we can recover (and display) selections from any viewpoint on the 3D shape in a few milliseconds.
In summary, we train a multiview-consistent 2D selection model and project selection similarities from multiple viewpoints to 3D using a simple, yet efficient, point cloud that can be queried from new cameras. Figure 1 provides an overview of our method.
@@ 3.1 Fine-tuning for 2D material selection
We re-purpose the recently introduced SAM2 model [49] to material selection in the video domain. SAM2 uses an efficient Vision Transformer (ViT) image encoder [52] to produce a per-frame image embedding, and infers a per-pixel object similarity value for each frame. The key novel component in SAM2 is the memory attention module, which conditions the current frame embedding on the embeddings of past and future frames in the sequence, allowing the model to reason both spatially and temporally. These embeddings are then combined with the encoded conditioning query (e.g., a click on a pixel) in the mask decoder, producing per-frame similarity masks. As we will show in our experiments, correctly fine-tuning this memory module is key to achieving multiview-consistent selection results.
@@
While our initial experiments confirmed SAM2’s good cross-view consistency, they also revealed a tendency to select object (sub-)parts instead of materials. We therefore fine-tune the model for the task of material selection. Specifically, we freeze the encoder throughout our fine-tuning to preserve the rich priors learned from millions of images and tune the remainder of the architecture (see Fig. 2). We find that training solely on image data for material selection (e.g., the Materialistic dataset [56]), performs reasonably well on clicked frames, but leads to a significant drop in cross-frame selection consistency, as shown in Fig. 3. We attribute this to the fact that on unseen frames, the model must infer the material selection from memory, but fine-tuning on images does not adapt the memory module since memory is never queried for a single image.
@@
However, for 3D selection, cross-frame consistency is particularly important. We therefore design an object-centric video dataset with material-segmentation annotation by randomly sampling objects, materials and environment maps, combining them into simple scenes containing one to a few objects. We allow the same materials to appear multiple times and in different locations within a scene, to clearly disambiguate material and object selection. We render 30 frames for each video using a random choice of four possible camera trajectories: zoom-in, zoom-out, spherical turntable and fly-over. Finally, to reduce the domain gap between natural and single-object images, we alpha-composit the environment map into the background (for additional dataset and training details, see Appendix A.1). We find that 500 videos are sufficient to adapt the model to the material-selection domain.
@@
Our new video material dataset with dense per-frame material annotations enables us to jointly fine-tune SAM2’s memory attention module and the mask decoder. This way, we maintain multiview-consistency and obtain significantly better inferred selections (right column in Fig. 3). We will release our dataset upon publication.
@@
Figure 3: Effects of fine-tuning on images vs. videos. Top row shows the clicked frame. Bottom row shows an unclicked frame for which the similarity map is inferred from the model’s memory.
@@ 3.2 Lifting 2D similarity to 3D
The initial camera, in which the click was performed, will serve to condition the memory module of our SAMa model, as it ensures that the material is not occluded. We then render RGB and depth images from multiple viewpoints; for each RGB image we use our model to estimate the selection, based on the user-provided click, given the other images as video context through the memory bank. This process yields a per-viewpoint map representing the similarity to the user-clicked material. We back-project these maps to 3D using the previously rendered depth images, to obtain a 3D similarity point cloud.
Our approach works on any 3D representation that can be rendered from a given viewpoint and queried for depth. For NeRFs and 3D Gaussians we use the training views, while for meshes we use spherical Fibonacci sampling for the camera positions (looking at the object’s center). To ensure good performance of our fine-tuned video model, we arrange those views along a smooth trajectory via greedy iterative camera sorting laid out in Appendix Alg. 1.
From our point cloud, we can reconstruct a continuous 3D similarity field via k-nearest neighbour (kNN) lookups. For novel views, we use FAISS [16] for performant large-scale, GPU-accelerated approximate nearest-neighbor queries [27] at the camera rays’ 3D hitpoints. We cache and reuse the acceleration structure built by the library; we need to rebuild it only when the selected material changes. With this approach, a new user selection from a novel viewpoint takes around 2 sec (including 0.5 sec for the structure construction), while querying the point cloud from a new viewpoint takes 10–20 ms at 512–1024p image resolution. Appendix A.2 provides additional details.
@@ 3.3 Refinement
We observe that the frame where the user clicks exhibits significantly higher selection noise. This is due to the memory module not being queried for this selection, meaning that the model does not have access to the information in the other frames. To improve selection quality on this frame, we simply duplicate it. The first copy is used for conditioning the selection without memory module, and the second copy is included with the other frames in the sequence, using the memory module. We show the effect of click-frame duplication in Fig. 4 and on the original SAM2 model in the supplemental material.
@@ 4.1 Datasets
We evaluate our method on three datasets: (1) the eight scenes in the NeRF dataset [42], (2) five real-world scenes from the MIPNeRF-360 dataset [2], and (3) twelve objects from our own object-centric dataset. For synthetic objects we render material-ID maps which provide ground-truth annotations per view. For real-world assets we hand-annotate five images per scene (examples in the supplemental).

### v2

@@ Abstract
Decomposing 3D assets into material parts is a common task for artists, yet remains a highly manual process. In this work, we introduce Select Any Material (SAMa), a material selection approach for in-the-wild objects in arbitrary 3D representations. Building on SAM2’s video prior, we construct a material-centric video dataset that extends it to the material domain. We propose an efficient way to lift the model’s 2D predictions to 3D by projecting each view into an intermediary 3D point cloud using depth. Nearest-neighbor lookups between any 3D representation and this similarity point cloud allow us to efficiently reconstruct accurate selection masks over objects’ surfaces that can be inspected from any view. Our method is multiview-consistent by design, alleviating the need for costly per-asset optimization, and performs optimization-free selection in seconds. SAMa outperforms several strong baselines in selection accuracy and multiview consistency and enables various compelling applications, such as replacing the diffuse-textured materials on a text-to-3D output with PBR materials or selecting and editing materials on NeRFs and 3DGS captures. Project page: https://mfischer-ucl.github.io/sama/.
@@ 1 Introduction
Understanding the materials around us is an extremely common task for humans, but remains challenging for machines. In this paper we focus on material selection in 3D.
Existing work on material understanding has mostly focused on the 2D image domain, addressing tasks like segmentation [65, 48, 67, 4], reconstruction [15, 16, 64, 30], generation [70, 60, 69] or, more recently, material selection [63]. Semantic segmentation and classification aim to separate areas into different predetermined classes, e.g., wood or plastic. This neither accounts for unforeseen materials nor the separation of two materials with different texture properties (e.g., two types of wood). This work, in contrast, targets selection, which aims at finding materials with the same appearance, and thus is more flexible as it can handle any material, regardless of class, and make intra-class distinctions such as two plastics with different appearance.
We follow the material definition of established works [63, 14, 26] and consider two materials similar only if they share the same texture and reflectance properties.
@@
Material selection becomes especially relevant in the light of recent generative 3D asset creation and image/text-to-3D workflows. Current methods either provide non-parametric implicit representations (e.g., Neural Radiance Fields) or unstructured output (as in triangle soups and baked textures produced by image/text-to-mesh methods [29, 45, 23, 52]), both of which are challenging to use for artists and downstream tasks. Material selection, in this context, has a wide range of downstream applications, e.g., enhancing the X-to-3D workflow with material masks, improving the editability of 3D reconstructions (e.g., through material replacement [75, 10]), or extracting areas of similar materials as a prior for inverse rendering [39].
@@
However, most models targeting material-related tasks, including selection, do not trivially extend to the 3D domain, as they are trained on 2D images and therefore have no incentive for producing multiview-consistent predictions [18]. Moreover, the 3D domain contains inherent challenges and ambiguities like self- or dis-occlusions and view-dependent effects, and requires accurate propagation of the model predictions into novel, unseen views. Recent research has therefore developed algorithms to address the problems from multiview-inconsistent predictions that arise when lifting 2D (object) selection to 3D, predominantly via pre-processing noise-consolidation steps such as feature-field distillation [36] or contrastive (similarity) learning [34], both of which are time-consuming (the former reports a 200k-step, the latter a 20-minute optimization) and must be repeated per asset.
@@
In this work, we close this gap between material selection in 2D and 3D by introducing Select Any Material (SAMa), an efficient and accurate material selection and segmentation method for 3D assets (see LABEL:fig:Teaser). Our first core insight is that we can draw parallels between video- and 3D-selection, since in both video and 3D, the selected elements have to be consistent across frames (or views), regardless of object and camera movement or differences in shading and occlusions. We thus propose to re-purpose SAM2’s recent progress in object selection across video [56] for materials. We achieve this by fine-tuning parts of the model for (video) material selection on a custom-made video dataset with dense per-pixel, per-frame annotations and show that using videos for fine-tuning is key to achieving high quality in 3D.
Our approach is inherently multiview-consistent thanks to its video training, which lets us avoid costly per-asset optimizations that lift the 2D signal to 3D and allows us resort to a simpler and more efficient strategy instead: for each pixel of each view, we project the 2D similarity information to 3D using the depth and simple inverse camera projections, thus constructing a 3D point cloud aligned with the 3D representation by construction. We achieve interactive visualization of our selection results by querying the nearest neighbor between the 3D representation and this similarity point cloud during rendering. As a result, our approach supports selection on any 3D representation that can be rendered to an image and queried for depth, such as meshes, Gaussian Splats and NeRFs. Additionally, it is fast: the similarly cloud reconstruction can be performed on-the-fly, while the selection visualization from novels views takes less than 10 milliseconds.
In summary, once our model is trained, it enables selection via multi-view renderings of an arbitrary 3D object in less than two seconds from the initial click.
@@
We evaluate our method on meshes, radiance fields and 3D Gaussians, in terms of selection quality and 3D consistency, and show that it improves significantly over existing work and several strong baselines. Finally, we demonstrate multiple applications such as object segmentation into material IDs and NeRF/Gaussian editing.
@@
  • Adaptation of a video-object-selection model to material selection on 3D shapes, by training on a novel rendered video dataset which will be released upon acceptance.
  • Fast and efficient 3D projections and queries, enabled by cross-frame consistent model predictions.
@@
Throughout this paper, we will show the user-provided input clicks with respect to which we select materials as black-rimmed circles, and the material similarity to these clicks in false colors, with blue and red indicating low and high, respectively.
@@ 2 Related work
Material selection.
Most prior works in material segmentation rely on hand-crafted features [58, 5, 47, 28, 55] or focus on images of flat surfaces [38, 30, 40, 51, 12]. Recently, Sharma et al. [63] proposed Materialistic, a model based on DINO-ViT [7] features, trained to predict the material similarity between a query pixel and all other pixels in a natural image. We find that Materialistic struggles with accurate material selection on 3D objects for two reasons: (1) it is trained for full-scene photographs, leading to limited selection precision on objects, and (2) its selections are not sufficiently consistent to be lifted to varying 3D views. For 3D segmentation, MatSeg3D [42] focuses on per-pixel material classification into a coarse set of predetermined categories (wood, metal, …), which prevents it from distinguishing between materials within the same category but different appearance (e.g. two different plastics).
Closely related, the Segment Anything Model (SAM) [35] uses a ViT trained to predict similarity between pixels. As its training data is object-selection specific and not material-aware, SAM requires many separate clicks to perform even moderately well on materials. The more recent SAM2 [56] also targets object selection, but introduces support for temporally coherent predictions across video frames. We find that neither SAM, SemanticSAM [41, 21], nor SAM2 perform well on material selection, except in the special case of an object made of a single material. However, once fine-tuned for materials, SAM2’s cross-frame selection consistency enables our fast selection lifting to 3D.
@@
Due to the scarcity of annotated 3D data and the increased computational complexity compared to 2D, many approaches attempt to lift 2D predictions from multiple views to a shared 3D representation.
@@ 3 Method
Our approach targets material selection on 3D representations. Given a 3D asset and a user click, we select all regions of the 3D asset sharing the material appearance of the clicked region. Existing methods focus mostly on selection in 2D images [63], and their extension to 3D is not trivial due to a lack of selection consistency across views.
Instead of enforcing 3D consistency through per-asset optimization and feature consolidation [36, 34, 25, 73], we note that recent video models [56, 68] show good cross-frame consistency. Since renders of a 3D object from a smooth camera trajectory are not markedly different from a video, we propose to adapt SAM2 [56] to material selection by fine-tuning it, including its memory bank components, on our material-specific video dataset. Once fine-tuned, the model consumes an image or video and outputs a per-frame floating-point map that encodes the similarity between the clicked pixel’s material and all other pixels.
However, fine-tuning is not sufficient for interactive selection in 3D. While it enables material selection from novel views with good consistency, each camera movement would require querying the model anew, with no guarantee of consistency in challenging cases on long frame sequences. To lift our selection to 3D, we consolidate the 2D similarity maps of a sparse set of key-frames into a 3D similarity point cloud. Combined with nearest-neighbor queries, we can recover (and display) selections from any viewpoint on the 3D shape in a few milliseconds. We show an overview of our method in Figure 1.
@@ 3.1 Fine-tuning for 2D material selection
We re-purpose the SAM2 model [56] to material selection in the video domain. SAM2 uses an efficient Vision Transformer (ViT) image encoder [59] to produce a per-frame image embedding, and infers a per-pixel object similarity value for each frame. The key novel component in SAM2 is the memory attention module, which conditions the current frame embedding on the embeddings of past and future frames in the sequence, allowing the model to reason both spatially and temporally. These embeddings are then combined with the encoded conditioning query (e.g., a click on a pixel) in the mask decoder, producing per-frame similarity masks. As we will show in our experiments, correctly fine-tuning this memory module is key to achieving multiview-consistent selection results.
@@
While our initial experiments confirmed SAM2’s good cross-view consistency, they also revealed a tendency to select based on semantic function (i.e. by parts) instead of materials. We therefore fine-tune the model for the task of material selection. Specifically, we freeze the encoder throughout our fine-tuning to preserve the rich priors learned from millions of images and tune the remainder of the architecture (see Fig. 3). We find that training solely on image data for material selection (e.g., the Materialistic dataset [63]), performs reasonably well on clicked frames, but leads to a significant drop in cross-frame selection consistency, as shown in Fig. 2. We attribute this to the fact that on unseen frames, the model must infer the material selection from memory, but fine-tuning on images does not adapt the memory module since it is never queried for a single image.
@@
However, for 3D selection, cross-frame consistency is particularly important. We therefore design an object-centric video dataset with material-segmentation annotation by randomly sampling objects, materials and environment maps, combining them into simple scenes containing one to a few objects. We allow the same materials to appear multiple times and in different locations within a scene, to clearly disambiguate material and object selection. We render 30 frames for each video using a random choice of four possible camera trajectories: zoom-in, zoom-out, spherical turntable and fly-over. Finally, to reduce the domain gap between natural and single-object images, we alpha-composite the environment map into the background (for additional dataset and training details, see Suppl. A.4. Since we freeze the pre-trained encoder to retrain its learned priors, we find that 500 videos are sufficient to adapt the model to the material-selection domain.
Figure 2: Fine-tuning on images creates artifacts on unseen frames, when selections are inferred from the model’s memory (right column).
@@
Our new video material dataset with dense per-frame material annotations lets us jointly fine-tune SAM2’s memory attention module and the mask decoder. This way, we maintain multiview consistency while significantly improving selections on unclicked frames (right column in Fig. 2). Our dataset and trained model are available on the project page and we show samples from the dataset in the supplemental.
@@ 3.2 Lifting 2D similarity to 3D
Given a click on the object, the initial camera, in which the click was performed, will serve to condition the memory module of our SAMa model, as it ensures that the material is not occluded. We then render RGB and depth images from multiple viewpoints; for each RGB image we use our model to estimate the selection, based on the user-provided click, given the other images as video context through the memory bank. This process yields a per-viewpoint map representing the similarity to the user-clicked material. We project these 2D similarity maps to 3D using the previously rendered depth images, to obtain a 3D similarity point cloud located near the object surface.
Our approach works on any 3D representation that can be rendered from a given viewpoint and queried for depth. For NeRFs and 3D Gaussians we use a subset of the training views, while for meshes we use spherical Fibonacci sampling of camera positions pointed at the object’s center. To ensure good performance of our fine-tuned video model, we arrange those views along a smooth trajectory via greedy iterative camera sorting described in Suppl. Algorithm 1.
For visualizing the 3D selection from novel views, we can reconstruct a continuous 3D similarity field via k-nearest neighbour (kNN) lookups into our previously constructed similarity point cloud. We use FAISS [17] for performant large-scale, GPU-accelerated approximate nearest-neighbor queries [31] at the camera rays’ 3D hitpoints. As our point cloud contains one point per observed pixel, it is very dense (see Suppl. Figure 9 for an example).
We cache and reuse the acceleration structure built by the library; we need to rebuild it only when the selected material changes. With this approach, a new user selection from a novel viewpoint takes around 2 sec (including 0.5 sec for the structure construction), while querying the point cloud from a new viewpoint takes 10–20 ms at 512–1024p image resolution. Suppl. Sec. A.2 provides additional details.
@@ 3.3 Refinement
We observe that the frame where the user clicks exhibits significantly higher selection noise. This is due to the fact that, by default, SAM2 does not query the memory module for this selection, meaning that the model does not have access to the information in the other frames. To improve selection quality on this frame, we simply duplicate it. The first copy is used for conditioning the selection without memory module, and the second copy is included with the other frames in the sequence, using the memory module. We show the benefit of click-frame duplication in Fig. 4 and on the original SAM2 model in the supplemental.
@@ 4.1 Datasets
We quantitatively evaluate our method on three datasets: (1) the eight scenes in the NeRF dataset [49], (2) five real-world scenes from MIPNeRF-360 [2], and (3) our object-centric dataset test split containing 12 objects. For synthetic assets we render material-ID maps which provide ground-truth annotations per view. For real-world assets we hand-annotate five images per scene (see Suppl. Fig. 2).
We show additional qualitative evaluations on real-world captures and photogrammetry output in the supplemental.